# Sharp Thresholds for Temporal Motifs and Doubling Time in Random Temporal Graphs

### v1

@@ 1 Introduction
A temporal graph is a pair $\mathcal{G}=(G,\lambda)$, where $G=(V,E)$ is an underlying (static) graph with $n$ denoting the number of vertices in $V$, and $\lambda:E\rightarrow 2^{\mathbb{R}_{+}}$ is a labeling function that assigns to every edge of $G$ a finite set of discrete time labels, each being a positive real number. $\mathcal{G}$ is a simple temporal graph if every edge has exactly one time label, i.e. the labeling function is $\lambda:E\rightarrow\mathbb{R}_{+}$.
@@
Motivated by the fact that, due to causality, information in temporal graphs can “flow” only along sequences of edges whose time-labels are increasing, the majority of temporal graph parameters and optimization problems that have been studied so far are based on the notion of temporal paths and other “path-related” notions, such as temporal analogues of distance, diameter, reachability, exploration, and centrality [AkridaGMS16, ErlebachHK15, MertziosMCS13, MichailS-TSP16, AkridaGMS-TOCS17, EnrightMMZ21, MolterRZ24, EnrightMM25, DogeasEKMM24, RymarMNN23, KlobasMMS25, KlobasMMNZ23]. Only relatively recently have attempts been made to define “non-path” temporal graph problems, such as temporal analogues of clique [viardClique, viardCliqueTCS, neidermeier, DavotEMM25], vertex cover [AkridaMSZ18, HammKMS22], coloring [MertziosMZ21, MarinoS22, AdamsonMS26, IbiapinaNRR25], matching [MertziosMNZZ20, BasteBR20, CioniD0S025], vertex separators [MaackMNR23, FluschnikMNRZ20, HarutyunyanKP23], and transitivity [MertziosMRSZ25].
@@
Continuing in this line of research on “non-path” temporal problems, we mainly focus in this paper on the existence of a given (arbitrary) temporal motif in a temporal graph. A temporal motif consists of a pair $(H,P)$, where $H$ is a fixed (static) graph and $P$ is a partial ordering over the edges of $H$. A temporal graph $\mathcal{G}=(G,\lambda)$ is said to contain a temporal motif $(H,P)$, if $H$ is contained in the footprint $G$, and the edges of $H$ have time labels which are ordered in agreement with the partial order $P$. The existence and frequency of temporal motifs is a question of both theoretical and practical importance. On the one hand, theoretically, temporal motifs stand as an intermediate object containing less temporal information than a temporal subgraph, but more temporal information than a footprint subgraph. In particular, they are well suited for capturing notions of causality and are less sensitive to the precise numerical timings of labels. On the other hand, temporal motifs are widely used in the practical analysis of temporal networks, for both identifying significant interaction patterns and determining the mechanics of network evolution. An especially common technique in this context is comparing the frequency of temporal motifs in a given empirical network with that of some appropriate (and typically adhoc) network. A strengthening of this concept, which we study, is the notion of a $\delta$-temporal motif (i.e. one where all of its constituent edge labels must lie within some window of length $\delta$). These, in turn, represent a middle ground between a temporal subgraph and a temporal motif. They have seen application both in fraud detection [FraudDetection] and as a tool for studying different interaction scales in empirical networks [DeltaTemporalMotifs].
@@
On the other hand, the vast majority of theoretical work on temporal graphs has focused on deterministic, worst-case inputs. While this has led to a relatively good understanding of the worst-case computational complexity of various problems on temporal graphs, this worst-case approach provides only little information about the vast majority of temporal graphs. This distinction was recently and dramatically demonstrated, by the discovery that almost all simple temporal graphs permit nearly optimal spanners [SharpThresholds], despite the existence of temporal graphs without sub-quadratic spanners [axiotis_et_al:LIPIcs.ICALP.2016.149]. Results such as these raise the general question: to what extent do these unique properties hold for random temporal graphs rather than only for carefully constructed ones?
@@ 1.1 Our Contribution
In this paper we focus on non-path problems on randomly generated temporal graphs. We expand on earlier work by [Cliques, SharpThresholds], by exploring the properties of two model variants (one continuous and one discrete model) for random temporal graphs, generalizing the Random Simple Temporal Graph Model (RSTG) [SharpThresholds, Cliques]. For ease of reading we defer the formal definition of the models (see Definitions˜2.3 and 2.4). Intuitively, in the continuous model we sample the time labels uniformly at random from $(0,1]$ for each edge, where the number of labels on each edge is taken from some discrete distribution. The discrete model is the natural discretizations of the continuous one, where the labels of each edge are chosen uniformly at random from a set $\{1,2,\ldots,T\}$. In our first result, we explore the existence thresholds for all fixed $\delta$-temporal motifs, in terms of the length $\delta$ of the permitted time window. We identify a behavior that is similar, but subtly different from the existence thresholds of fixed static graphs in the Erdős-Rényi random graph model. In particular, the threshold is determined not by the reciprocal of the density of the densest subgraph, but instead by the following quantity which we call the sparsity.
@@
This seemingly minor change is sufficient to cause qualitatively different threshold behavior from the static case. For example, motifs built from cycles of different sizes have asymptotically distinct existence thresholds, unlike their static counterparts [frieze2015introduction]. Another consequence of this result is that the asymptotic thresholds are independent of the edge ordering of a given motif. This implies, for example, that time respecting paths of a given length have the same asymptotic threshold as non-time respecting paths of the same length. Furthermore, we extend this behavior beyond fixed motifs to consider the size of the largest temporal motif with a clique footprint, and obtain tight bounds in the case where the number of labels associated with each edge is concentrated on at most two values.
Our second main result considers a different celebrated property of static random graphs: their higher order connectivity. Since the existence of a good analogue for expansion in temporal graphs remains an open problem, we use the maximum doubling time of the reachability ball centered on small sets of vertices as a natural proxy. In our continuous model, when the number of labels per edge is fixed, we provide sharp upper and lower bounds on the maximum doubling time. As a direct consequence, we immediately obtain a sharp threshold for finite maximum doubling time in the Random Simple Temporal Graph (RSTG) model of [SharpThresholds]. Interestingly, this corresponds precisely to the threshold for the existence of temporal source and sink vertices in that model. In contrast, the doubling time of any specific large set of vertices corresponds to the threshold for both point to point reachability and the existence of a giant component in the RSTG.
@@ 2 Model and Definitions
As a natural generalization of an important tool for the analysis of static graphs, some notion of “temporal motifs” have been studied in a variety of contexts and with a variety of definitions (see [sariyuce2025powerful] for an excellent survey). Particular points of distinction relevant to this work include whether an occurrence of a temporal motif must represent an induced subgraph (see [DeltaTemporalMotifs] and [AnalyticalModels] for a difference in opinion) and whether the associated order must be partial [kovanen2011temporal], or total [DeltaTemporalMotifs, AnalyticalModels, SamplingMethodsForCountingTemporalMotifs, 10.1093/bioinformatics/btv227]. We make use of the following definition for (simple) $\delta$-temporal motifs.
@@
  • $\mathcal{I}$ is a simple temporal subgraph of $G$: $V^{\prime}\subseteq V$, $E^{\prime}\subseteq E[V^{\prime}]$ and $\forall e\in E^{\prime}:\lambda^{\prime}(e)\in\lambda(e)$
  • $\Phi$ is an isomorphism of $H$ and realizes the order over the labels: $\Phi$ is an isomorphism of $H$, $\forall uv,wx\in E_{H}:uv\prec_{P}wx\implies\lambda^{\prime}(\Phi(u)\Phi(v))<\lambda^{\prime}(\Phi(w)\Phi(x))$  [^1] Definition 2.1 is given under the assumption that $P$ is a strict partial order. Alternatively a $\delta$-temporal motif could be defined with respect to a non-strict partial order $P$, that is, $uv\prec_{P}wx\implies\lambda^{\prime}(\Phi(u)\Phi(v))\leq\lambda^{\prime}(\Phi(w)\Phi(x))$. With this alternative definition, all results of the paper would remain the same.
@@
We define the label multiplicity distribution $\psi$ to be any non-negative discrete distribution, such that $\psi$ is non-zero with positive probability and both the first and second moments exist.
For a random variable $Q\sim\psi$, we will take $r=\mathbb{E}[Q]$ and $r_{2}=\mathbb{E}[Q^{2}]$. Furthermore, we take $q=\mathbb{P}[Q>0]$. We can now define both our continuous and discrete models.
@@
    The doubling time $\textbf{Double}(\mathcal{G})=\max_{S\subset V:|S|\leq\frac{|V|}{2}}\text{Double}_{\mathcal{G}}(S)$ of a temporal graph to be the maximum doubling time over all sets of vertices
@@ 3 Results
Let $H=(V_{H},E_{H})$ be a static graph, $P=(E_{H},\prec)$ be any partial order over the edges of $H$ and $\rho_{H}=\min_{I\subseteq H}\frac{V_{I}}{E_{I}-1}$. Suppose that there exists $n_{0}>0$ such that $\min_{n\geq n_{0}}\beta(n)\geq\text{height}(P)$. Then, a random temporal graph drawn from $\Gamma_{[n]}(\psi,T(n))$
@@
Let $C_{i}$ be the cycle on $i$ vertices and $P_{i}$ be any partial order over its edge sets. A random temporal graph drawn from $\Gamma_{[n]}(\psi)$:
@@
Let $K_{i}$ be the clique on $i$ vertices and $P^{\emptyset}_{i}$ be the partial order over its edges such that they are all mutually incomparable. Further, let $\delta>0$ be a constant, $\mathcal{G}=(V,E,\lambda)$ be a sample from $\Gamma_{[n]}(\psi)$ and define $w_{\delta}(\mathcal{G})$ to be the maximum value such that $\mathcal{G}$ contains $(H_{w_{\delta}(\mathcal{G})},P^{\emptyset}_{w_{\delta}(\mathcal{G})})$ as a $\delta$-temporal motif. Then for any $\epsilon>0$, the following holds with high probability,
@@
$\mathbb{P}\left[\frac{(2-\alpha)\log{n}}{r(n-1)}\leq\mathbf{Double}(\mathcal{G})\leq\frac{(2+\alpha)\log{n}}{r(n-1)}\right]=1-o(1).$
@@ 4.1 Technical Preliminaries
Before we present our proofs, we will need a few preliminary results and definitions. Our first preliminary is a short result concerning the probability of a given collection of labels obeying the partial order. Here it is important to distinguish between the continuous case, where labels are totally ordered almost surely, and the discrete case, where multiple labels may share the same time stamp.
@@
We will also need the following notion of automorphism for temporal motifs. In particular, we require that the mapping over the edges produced by the automorphism of the vertices corresponds to an automorphism of the associated partial order. The number of such automorphisms will play a crucial role in preventing over-counting of the number of occurrences of temporal motifs when multiple occurrences share the same labels.
@@
We define $\mathcal{S}$ to be the set of all occurrences of $H$ in the footprint graph $G=(V,E)$ and for any $S=(V_{S},E_{S})$ we define $\mathcal{T}_{S}$ to be the set of all mappings from $E_{S}\to\mathbb{N}_{0}$. Furthermore, for mathematical convenience we shall treat $\lambda(e)$ as a tuple ordered by the order of sampling rather than a multi-set and we shall assume that for each edge the value of all labels are sampled first, followed by how many will actually appear.
For a random sample $\mathcal{G}$ from $\Gamma_{[n]}(\psi)$, $S\in\mathcal{S}$ and $\tau\in\mathcal{T}_{S}$, we define $C^{(\delta)}_{S,\tau}$ to be the indicator random variable for the event that $(S,\tau)$ describes a $\delta$-temporal $(H,P)$ motif of $\mathcal{G}$. Additionally, $\mathbf{C}_{\delta}=\sum_{S\in\mathcal{S},\tau\in\mathcal{T}_{S}}C^{(\delta)}_{S,\tau}$ then corresponds to the total number of appearances of $(H,P)$ as a $\delta$-temporal motif of $\mathcal{G}$. Analogously, for a random sample $\mathcal{H}$ from $\Gamma_{[n]}(\psi,T)$, which we obtain by taking the $T$-discretization of the same sample of $\Gamma_{[n]}(\psi)$, we define $F^{(\beta),T}_{S,\tau}$ to be the indicator for the event that $(S,\tau)$ describes a $\beta$-temporal $(H,P)$ motif of $\mathcal{H}$. Finally, $\mathbf{F}_{(\beta),T}=\sum_{S\in\mathcal{S},\tau\in\mathcal{T}_{S}}F^{(\beta),T}_{S,\tau}$ corresponds to the total number of appearances of $(H,P)$ as a $\beta$-temporal motif of $\mathcal{H}$.
We now make the following observation relating the two models. {observation}[] For $\mathcal{G}=(V,E,\lambda)$, $\beta,T>1$, $H=(V_{H},E_{H})$ any static graph, $P=(E_{H},\prec)$ a partial order over the edges of $H$, and for every $S\in\mathcal{S}$, $\tau\in\mathcal{T}_{S}$, we have that $\{F^{(\beta),T}_{S,\tau}=1\}\subseteq\{C^{(\frac{\beta}{T})}_{S,\tau}=1\}.$ Therefore, with certainty, $\mathbf{F}_{(\beta),T}\leq\mathbf{C}_{\frac{\beta}{T}}.$

### v2

@@ 1 Introduction
A temporal graph is a pair $\mathcal{G}=(G,\lambda)$, where $G=(V,E)$ is an underlying (static) graph with $n$ denoting the number of vertices in $V$, and $\lambda:E\rightarrow 2^{\mathbb{R}_{+}}$ is a labeling function that assigns to every edge of $G$ a finite set of time labels, each being a positive real number. $\mathcal{G}$ is a simple temporal graph if every edge has exactly one time label, i.e. the labeling function is $\lambda:E\rightarrow\mathbb{R}_{+}$.
@@
Temporal graphs are traditionally used to model information flow which, due to causality, can only pass along a sequence of edges with increasing time labels. Motivated by this, the majority of work concerning temporal graphs is based on the notion of temporal paths and other “path-related” notions, such as temporal analogues of distance, diameter, reachability, exploration, and centrality [AkridaGMS16, ErlebachHK15, MertziosMCS13, MichailS-TSP16, AkridaGMS-TOCS17, EnrightMMZ21, MolterRZ24, EnrightMM25, DogeasEKMM24, RymarMNN23, KlobasMMS25, KlobasMMNZ23]. Only relatively recently have attempts been made to define “non-path” temporal graph problems, such as temporal analogues of clique [viardClique, viardCliqueTCS, neidermeier, DavotEMM25], vertex cover [AkridaMSZ18, HammKMS22], coloring [MertziosMZ21, MarinoS22, AdamsonMS26, IbiapinaNRR25], matching [MertziosMNZZ20, BasteBR20, CioniD0S025], vertex separators [MaackMNR23, FluschnikMNRZ20, HarutyunyanKP23], and transitivity [MertziosMRSZ25].
@@
Continuing in this line of research on “non-path” temporal problems, we mainly focus in this paper on the existence of a given (arbitrary) temporal motif in a temporal graph. A temporal motif consists of a pair $(H,P)$, where $H$ is a fixed (static) graph and $P$ is a partial ordering over the edges of $H$. A temporal graph $\mathcal{G}=(G,\lambda)$ is said to contain a temporal motif $(H,P)$, if $H$ is contained in the footprint $G$, and the edges of $H$ have time labels which are ordered in agreement with the partial order $P$. The existence and frequency of temporal motifs is a question of both theoretical and practical importance. On the one hand, theoretically, temporal motifs stand as an intermediate object containing less temporal information than a temporal subgraph, but more temporal information than a footprint subgraph. In particular, they are well suited for capturing notions of causality and are less sensitive to the precise numerical timings of labels. On the other hand, temporal motifs are widely used in the practical analysis of temporal networks, for both identifying significant interaction patterns and determining the mechanics of network evolution [sariyuce2025powerful]. An especially common technique in this context is comparing the frequency of temporal motifs in a given empirical network with that of some appropriate (and frequently ad hoc) network[sariyuce2025powerful, doi:10.1137/19M1242252, kovanen2011temporal]. A strengthening of this concept, which we study, is the notion of a $\delta$-temporal motif (i.e. one where all of its constituent edge labels must lie within some window of length $\delta$). These, in turn, represent a middle ground between a temporal subgraph and a temporal motif. They have seen application both in fraud detection [FraudDetection] and as a tool for studying different interaction scales in empirical networks [DeltaTemporalMotifs].
@@
Additionally, the vast majority of theoretical work has focused on deterministic worst-case temporal graphs. While this has led to a relatively good understanding of the worst-case temporal graph properties, the approach provides little information about the vast majority of temporal graphs. This distinction was recently and dramatically demonstrated, by the discovery that almost all simple temporal graphs permit nearly optimal spanners [SharpThresholds], despite the existence of temporal graphs without sub-quadratic spanners [axiotis_et_al:LIPIcs.ICALP.2016.149]. Results such as these raise the general question: to what extent do such unique properties hold for random temporal graphs rather than only for carefully constructed ones?
@@ 1.1 Our Contribution
In this paper we focus on non-path problems on randomly generated temporal graphs. We expand on earlier work by [Cliques, SharpThresholds], by exploring the properties of two model variants (one continuous and one discrete model) for random temporal graphs, generalizing the Random Simple Temporal Graph Model (RSTG) [SharpThresholds, Cliques]. For ease of reading we defer the formal definition of the models (see Definitions˜2.3 and 2.4). Intuitively, in the continuous model we sample the time labels uniformly at random from $(0,1]$ for each edge, where the number of labels on each edge is sampled independently for each each edge from some discrete distribution. The discrete model is the natural discretization of the continuous one, where the labels of each edge are chosen uniformly at random from a set $\{1,2,\ldots,T\}$. In our first result, we explore the existence thresholds for all fixed $\delta$-temporal motifs, in terms of the length $\delta$ of the permitted time window. We identify a behavior that is similar, but subtly different from the existence thresholds of fixed static graphs in the Erdős-Rényi random graph model. In particular, the threshold is determined not by the reciprocal of the density of the densest subgraph, but instead by the following quantity which we call the sparsity.
@@
This seemingly minor change is sufficient to cause qualitatively different threshold behavior from the static case. For example, motifs built from cycles of different sizes have asymptotically distinct existence thresholds, unlike their static counterparts [frieze2015introduction]. Another consequence of this result is that the asymptotic thresholds are independent of the edge ordering of a given motif. This implies, for example, that time-respecting paths of a given length have the same asymptotic threshold as non-time-respecting paths of the same length. Furthermore, we extend this behavior beyond fixed motifs to consider the size of the largest temporal motif with a clique footprint, and obtain tight bounds in the case where the distribution of labels associated with each edge is supported on only one non-zero value.
Our second main result considers a different celebrated property of static random graphs: their higher order connectivity. Since the existence of a good analogue for expansion in temporal graphs remains an open problem, we use the maximum doubling time of the reachability ball centered on small sets of vertices as a natural proxy. In our continuous model, when the number of labels per edge is fixed, we provide sharp upper and lower bounds on the maximum doubling time. As a direct consequence, we immediately obtain a sharp threshold for finite maximum doubling time in the Random Simple Temporal Graph (RSTG) model of [SharpThresholds]. Interestingly, this corresponds precisely to the threshold for the existence of temporal source and sink vertices in that model. In contrast, the doubling time of any specific large set of vertices corresponds to the threshold for both point-to-point reachability and the existence of a giant component in the RSTG.
@@ 2 Model and Definitions
As a natural generalization of an important tool for the analysis of static graphs, some notion of a “temporal motif” has been studied in a variety of contexts and with a variety of definitions (see [sariyuce2025powerful] for a survey). Particular points of distinction relevant to this work include whether an occurrence of a temporal motif must represent an induced subgraph (see [DeltaTemporalMotifs] and [AnalyticalModels] for a difference in opinion) and whether the associated order must be partial [kovanen2011temporal], or total [DeltaTemporalMotifs, AnalyticalModels, SamplingMethodsForCountingTemporalMotifs, 10.1093/bioinformatics/btv227]. We make use of the following definition for (simple) $\delta$-temporal motifs.
@@
  • $\mathcal{I}$ is a simple temporal subgraph of $\mathcal{G}$: $V^{\prime}\subseteq V$, $E^{\prime}\subseteq E[V^{\prime}]$ and $\forall e\in E^{\prime}:\lambda^{\prime}(e)\in\lambda(e)$
  • $\Phi$ is an isomorphism of $H$ and realizes the order over the labels: $\Phi$ is an isomorphism of $H$, $\forall uv,wx\in E_{H}:uv\prec_{P}wx\implies\lambda^{\prime}(\Phi(u)\Phi(v))<\lambda^{\prime}(\Phi(w)\Phi(x))$.  [^1] Definition 2.1 is given under the assumption that $P$ is a strict partial order. Alternatively a $\delta$-temporal motif could be defined with respect to a non-strict partial order $P$, that is, $uv\prec_{P}wx\implies\lambda^{\prime}(\Phi(u)\Phi(v))\leq\lambda^{\prime}(\Phi(w)\Phi(x))$. With this alternative definition, all results of the paper would remain the same.
@@
We define the label multiplicity distribution $\psi$ to be any non-negative discrete distribution with defined positive and finite first and second moments.
For a random variable $Q\sim\psi$, we denote $r=\mathbb{E}[Q]$ and $r_{2}=\mathbb{E}[Q^{2}]$. Furthermore, we denote $q=\mathbb{P}[Q>0]$. We can now define both our continuous and discrete models.
@@
    The doubling time $\textbf{Double}(\mathcal{G})=\max_{S\subset V:|S|\leq\frac{|V|}{2}}\text{Double}_{\mathcal{G}}(S)$ of a temporal graph to be the maximum doubling time over all small sets of vertices
@@ 3 Results
Let $H=(V_{H},E_{H})$ be a static graph, $P=(E_{H},\prec)$ be any partial order over the edges of $H$ and $\rho_{H}=\min_{I\subseteq H}\frac{V_{I}}{E_{I}-1}$. Suppose that there exists $n_{0}>0$ such that $\min_{n\geq n_{0}}\beta(n)\geq\text{height}(P)$. Then, for $T(n)\geq\beta(n)$, a random temporal graph drawn from $\Gamma_{[n]}(\psi,T(n))$
@@
Let $C_{i}$ be the cycle on $i$ vertices and $P_{i}$ be any partial order over its edge set. A random temporal graph drawn from $\Gamma_{[n]}(\psi)$:
@@
Let $K_{i}$ be the clique on $i$ vertices and $P^{\emptyset}_{i}$ be the partial order over its edges such that they are all mutually incomparable. Further, let $0<\delta<1$ be a constant, $\mathcal{G}=(V,E,\lambda)$ be a sample from $\Gamma_{[n]}(\psi)$ and define $w_{\delta}(\mathcal{G})$ to be the maximum value such that $\mathcal{G}$ contains $(H_{w_{\delta}(\mathcal{G})},P^{\emptyset}_{w_{\delta}(\mathcal{G})})$ as a $\delta$-temporal motif. Then for any $\epsilon>0$, the following holds with high probability,
@@
$\mathbb{P}\left[\frac{(2-\alpha)\log{n}}{rn}\leq\mathbf{Double}(\mathcal{G})\leq\frac{(2+\alpha)\log{n}}{rn}\right]=1-o(1).$
@@ 4.1 Technical Preliminaries
Before we present our techniques, we will need a few preliminary results and definitions. Our first preliminary is a short definition required for defining the probability of a given collection of labels obeying the partial order. Here it is important to distinguish between the continuous case, where labels are totally ordered almost surely, and the discrete case, where multiple labels may share the same time stamp.
@@
We will also need the following notion of automorphism for temporal motifs. In particular, we require that the mapping over the edges produced by the automorphism of the vertices corresponds to an automorphism of the associated partial order. The number of such automorphisms will play a crucial role in preventing the over-counting of the number of occurrences of temporal motifs when multiple occurrences share the same labels.
@@
We define $\mathcal{S}$ to be the set of all occurrences of $H$ in the footprint graph $G=(V,E)$ and for any $S=(V_{S},E_{S})$ we define $\mathcal{T}_{S}$ to be the set of all mappings from $E_{S}\to\mathbb{N}_{0}$. Furthermore, for mathematical convenience we shall treat $\lambda(e)$ as a tuple ordered by the order of sampling rather than a multi-set and we shall assume that for each edge the value of all labels are sampled first, followed by how many will actually appear. Then, for a random sample $\mathcal{G}$ from $\Gamma_{[n]}(\psi)$, $S\in\mathcal{S}$ and $\tau\in\mathcal{T}_{S}$, we define $C^{(\delta)}_{S,\tau}$ to be the indicator random variable for the event that $(S,\tau)$ describes a $\delta$-temporal $(H,P)$ motif of $\mathcal{G}$. Additionally, $\mathbf{C}_{\delta}=\sum_{S\in\mathcal{S},\tau\in\mathcal{T}_{S}}C^{(\delta)}_{S,\tau}$ corresponds to the total number of appearances of $(H,P)$ as a $\delta$-temporal motif of $\mathcal{G}$. Analogously, for a random sample $\mathcal{H}$ from $\Gamma_{[n]}(\psi,T)$, which we obtain by taking the $T$-discretization of $\mathcal{G}$, we define $F^{(\beta),T}_{S,\tau}$ to be the indicator for the event that $(S,\tau)$ describes a $\beta$-temporal $(H,P)$ motif of $\mathcal{H}$. Finally, $\mathbf{F}_{(\beta),T}=\sum_{S\in\mathcal{S},\tau\in\mathcal{T}_{S}}F^{(\beta),T}_{S,\tau}$ corresponds to the total number of appearances of $(H,P)$ as a $\beta$-temporal motif of $\mathcal{H}$.
We now make the following observation relating the two models. {observation}[] For $\mathcal{G}=(V,E,\lambda)$, $T>\beta\geq 1$, $H=(V_{H},E_{H})$ any static graph, $P=(E_{H},\prec)$ a partial order over the edges of $H$, and for every $S\in\mathcal{S}$, $\tau\in\mathcal{T}_{S}$, we have that $\{F^{(\beta),T}_{S,\tau}=1\}\subseteq\{C^{(\frac{\beta}{T})}_{S,\tau}=1\}.$ Therefore, with certainty, $\mathbf{F}_{(\beta),T}\leq\mathbf{C}_{\frac{\beta}{T}}.$